# Exact Fisher zeros and thermofield dynamics across a quantum critical point

### v2

@@ I Discussion on the Suzuki solution
I.1 Numerical check of original Suzuki solution
The original Suzuki solution [1] of the partition function is
$\begin{split}Z_{\rm{su}}&=\frac{1}{2}\left[\prod_{k=1}^{L}2\cosh(\beta\epsilon_{2k})+\prod_{k=1}^{L}2\sinh(\beta\epsilon_{2k})\right.\\ &\left.+\prod_{k=1}^{L}2\cosh(\beta\epsilon_{2k-1})+\prod_{k=1}^{L}2\sinh(\beta\epsilon_{2k-1})\right].\end{split}$ (S1)
where the Kramers-Wannier (KW) [2] is fully manifested; specifically, Eq. S1 remains invariant under the transformation $g\rightarrow 1/g$, $\beta\rightarrow\beta g$. Consequently, the Fisher zero results for $g>1$ can be easily derived by mapping to the dual space ($g<1$). Therefore, zero lines will reappear for $g>1$ to characterize the dual spin flip excitations, similar to those for domain walls. Figure S1(a) shows the results of Fisher zeros at $g=1.5$ for $L=16$ and in the thermodynamic limit, according to the Suzuki solution and the exact formula given in Ref. [3], thereby illustrating the KW duality behavior between spin flips and domain walls on the two sides of the QCP in the complex $\beta$ plane.
To verify whether the physical picture at $g>1$ obtained from the Suzuki solution is correct, we employ two precise numerical methods, ED and tensor network calculations, to obtain Fisher zeros solutions for the same small system with $L=16$. In the tensor network calculations, we first construct a 2D lattice of size $L\times n$ from the 1D system by using Trotter decomposition with $\beta=\tau n$, then solve the partition function using the higher-order tensor renormalization group (HOTRG) method [4], an efficient tensor network algorithm capable of handling sign problems in the complex $\beta$ plane [5]. In our HOTRG calculations, $n=$1000, and the bond dimension $D_{b}$ describing the entanglement properties is set to 32. We find that the ED and the HOTRG calculations provide consistent results [Fig. S1(b)], but they are entirely inconsistent with the picture from the Suzuki solution with the same system size $L=16$. In the latter two numerical results, there is no reappearance of Fisher zero lines in the complex $\beta$ plane; instead, only localized Fisher zero structures are found [Fig. S1(b)], indicating the presence of closed zero curves in the thermodynamic limit. This implies that a new description is needed for spin flip excitations in the quantum disordered phase on the complex $\beta$ plane, as the Fisher zero lines strictly adhering to KW duality according to the Suzuki solution will no longer apply.
@@ I.1 Numerical check of original Suzuki solution
I.2 Modified Suzuki solution at quantum disordered side
In this section, we discuss the reasons for the discrepancy between results from the Suzuki solution and those from the unbiased numerical calculations. Prior to deriving Eq. S1, anisotropic Onsager/Kaufman results are used, with the Trotter step $n$ tends to infinity:
@@ I.2 Modified Suzuki solution at quantum disordered side
Differing from the ordered side, where the system has a degeneracy of two, on the quantum disordered side, the system has a degeneracy of only one. This can be addressed by selecting opposite signs for $\gamma_{2L}$ on different sides, thereby ensuring that $\lambda_{-}/\lambda_{+}=1$ for $g<g_{c}$ and $\lambda_{-}/\lambda_{+}<1$ for $g>g_{c}$. Therefore, at $n\rightarrow\infty$, $\gamma_{k}=2\beta\epsilon_{k}/n$ , except for $\gamma_{2L}=2\beta\epsilon_{0}/n$ at $g<g_{c}$ and $\gamma_{2L}=-2\beta\epsilon_{0}/n$ at $g>g_{c}$, respectively. Actually, substituting the specific form of $\epsilon_{0}$ into $\gamma_{2L}$ makes it easier to understand. In this case, $\gamma_{2L}=2\beta(g_{c}-g)/n$. From an algebraic perspective, the old form $\gamma_{2L}=2\beta|g_{c}-g|/n$ is symmetric at the QCP but not differentiable, leading to the reappearance of zero lines at $g>g_{c}$ and creating a fake illusion of KW duality on the complex $\beta$ plane. Then the exact Suzuki solution at $g>g_{c}$ need to modify as
@@
II fisher zeros solution and velocity of zeros in the thermodynamic limit
$\tilde{S}=\int_{0}^{\pi}\log|\tanh(\beta\epsilon_{q})|^{2}dq=0$ (S6)
$v_{g}=\frac{\partial\beta}{\partial g}=-\left(\frac{\partial\tilde{S}}{\partial g}\right)_{\beta}/\left(\frac{\partial\tilde{S}}{\partial\beta}\right)_{g}$ (S7)
III Scaling of the Zeros
Figure S4: The oscillatory behavior of $S$ at different $L$ values not only exhibits the same periodicity but also shows nearly perfect overlap, demonstrating self-similar behavior. Examples with (a) $n=2$ and (b) $n=3$.
IV self similarity of $S$
Self-similarity is a fundamental characteristic of quantum critical points, reflecting the scale-invariant nature of quantum fluctuations at these points. The relationship $S(\beta_{r},t)_{L}\approx S(n\beta_{r},nt)_{nL}$ at $g_{c}=1$ provides a classic example of how systems at different scales can exhibit similar behavior. This property can be used for RG analysis, a powerful tool in studying critical phenomena. Figure S4 illustrates the self-similar behavior of $S$ using $n=2$ and $n=3$ as examples. The oscillatory behavior of S at different L values not only exhibits the same periodicity but also shows nearly perfect overlap, demonstrating the self-similar nature. The self-similarity implies that the critical exponents, which characterize the behavior of physical quantities near the QCP, are scale-invariant.

### v3

@@ I Discussion on the Suzuki solution
I.1 Comparing the original Suzuki solution to HOTRG and ED
The original Suzuki solution [1] of the partition function of the one-dimensional transverse field Ising Model (1DTFIM) is
$\begin{split}Z_{\rm{su}}&=\frac{1}{2}\left[\prod_{k=1}^{L}2\cosh(\beta\epsilon_{2k})+\prod_{k=1}^{L}2\sinh(\beta\epsilon_{2k})\right.\\ &\left.+\prod_{k=1}^{L}2\cosh(\beta\epsilon_{2k-1})+\prod_{k=1}^{L}2\sinh(\beta\epsilon_{2k-1})\right],\end{split}$ (S1)
where $\epsilon_{k}=[1+g^{2}-2g\cos(\pi k/L)]^{1/2}$ and the Kramers-Wannier (KW) [2] is fully manifested. Specifically, Eq. S1 remains invariant under the transformation $g\rightarrow 1/g$, $\beta\rightarrow\beta g$. Consequently, the Fisher zero results for $g>1$ can be easily derived by mapping to the dual space ($g<1$). Therefore, zero lines will reappear for $g>1$ to characterize the dual spin flip excitations, similar to those for domain walls. Figure S1(a) shows the results of Fisher zeros at $g=1.5$ for $L=16$ and in the thermodynamic limit, according to the Suzuki solution and the exact formula given in Ref. [3], thereby illustrating the KW duality behavior between spin flips and domain walls on the two sides of the QCP in the complex $\beta$ plane.
To verify whether the physical picture at $g>1$ obtained from the Suzuki solution is correct, we employ two precise numerical methods, exact diagonalization (ED) and tensor network calculations, to obtain Fisher zeros solutions for the same small system with $L=16$. In the tensor network calculations, we first construct a 2D lattice of size $L\times n$ from the 1D system by using Trotter decomposition with $\beta=\tau n$, then solve the partition function using the higher-order tensor renormalization group (HOTRG) method [4], an efficient tensor network algorithm capable of handling sign problems in the complex $\beta$ plane [5]. In our HOTRG calculations, $n=$1000, and the bond dimension $D_{b}$ describing the entanglement properties is set to 32. We find that the ED and the HOTRG calculations provide consistent results [Fig. S1(b)], but they are entirely inconsistent with the picture from the Suzuki solution with the same system size $L=16$. In the latter two numerical results, there is no reappearance of Fisher zero lines in the complex $\beta$ plane; instead, only localized Fisher zero structures are found [Fig. S1(b)], indicating the presence of closed zero curves in the thermodynamic limit. This implies that a new description is needed for spin flip excitations in the quantum disordered phase on the complex $\beta$ plane, as the Fisher zero lines strictly adhering to KW duality according to the Suzuki solution will no longer apply.
@@ I.1 Comparing the original Suzuki solution to HOTRG and ED
I.2 Modified Suzuki solution on quantum disordered side
We then discuss the reasons for the discrepancy between results from the Suzuki solution and those from the unbiased numerical calculations. Prior to deriving Eq. S1, anisotropic Onsager/Kaufman results are used, with the Trotter step $n$ tends to infinity:
@@ I.2 Modified Suzuki solution on quantum disordered side
[…]
@@
II Structure of Fisher zeros
II.1 Finding Fisher zeros from the analytical solution of $Z$ for the 1DTFIM
With the form of $Z$ determined, we now sketch how Fisher zeros are located and visualized. To analytically continue $Z$, we simply allow $\beta$ to be complex, $\beta=\beta_{r}+i\beta_{i}$, in Eq. (2) in the main text. On the complex $\beta$ plane, $Z(\beta)$ is not a polynomial but an entire function of $\beta$. Following Fisher [7], we can factorize $Z$ into $Z(\beta)=e^{h(\beta)}\prod_{j}(\beta-z_{j})$ where $h$ is an entire function, and the multiplicity of $z_{j}$, the $j$-th root of $Z=0$, is assumed to be 1. Define the free energy (up to a factor of $\beta$) as $f=-\ln Z=-h-f_{na}$, where the non-analytical part $f_{na}=\sum_{j}\ln(\beta-z_{j})$ develops singularities at $\beta=z_{j}$ and is completely determined by the Fisher zeros $\{z_{j}\}$. Let $f_{na}=\phi+i\psi$, then real part $\phi=\sum_{i}\ln|\beta-z_{j}|$ can be interpreted as the electric potential due to a set of point charges at $\{z_{j}\}$ [8]. The electric field lines stem out of these Fisher charges and correspond to the contour line of the imaginary part $\psi$ [8]. A branch cut of $f$ is developed at each Fisher zero. Thus the contour plot of the real and imaginary part of $f$ can serve as an intuitive way to locate the Fisher zeros, matching the exact solution (Fig. S3). Note that in Fig. S3, due to the mirror symmetry of Fisher zeros with respect to the $\beta_{i}$ axis, only the Fisher zeros at the first quadrant is displayed. In the thermodynamic limit, if the Fisher zeros come close to each other, the charge density $\rho(\beta)=\sum_{j}\delta(\beta-z_{j})$ may become a continuous function.
Figure S3: Locating Fisher zeros. (a) The contour lines of ${\rm Re}f$. The white lines indicate the branch cuts, which terminate the Fisher zeros at the starting points ($L=32$,$g=1.8$). (b) Zeros of the same system determined by the intersections of ${\rm Re}(Z)=0$ and ${\rm Im}(Z)=0$ from the exact solution.
II.2 Velocity of zeros of the 1DTFIM in the thermodynamic limit
$\tilde{S}=\int_{0}^{\pi}\log|\tanh(\beta\epsilon_{q})|^{2}dq=0,$ (S6)
$v_{g}=\frac{\partial\beta}{\partial g}=-\left(\frac{\partial\tilde{S}}{\partial g}\right)_{\beta}/\left(\frac{\partial\tilde{S}}{\partial\beta}\right)_{g},$ (S7)
II.3 Fisher zeros in other spin models
We first consider the ANNNI model
$H=-\sum_{i=1}^{L}\left(\sigma_{i}^{z}\sigma_{i+1}^{z}-\kappa\sigma_{i}^{z}\sigma_{i+2}^{z}+g\sigma_{i}^{x}\right),$ (S8)
where $\kappa$ is the next-nearest-neighbor antiferromagnetic coupling and $g$ is the transverse field. In the case where $\kappa$ and $g$ are relatively small, the excitations can still be described by domain walls. Results of the Fisher zeros using ED and tensor network methods consistently show that the open line-shaped zeros [Fig. S4(a)], which describe domain wall excitations in the 1DTFIM, still exist. The introduction of $\kappa$ only causes this line to tilt.
We further consider the XY model
$H=\sum_{i=1}^{L}[(1+\gamma)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\gamma)\sigma_{i}^{y}\sigma_{i+1}^{y}],$ (S9)
where at $\gamma=0$, it transforms into the gapless XX model. We choose a small $g$ at $g=0.01$ to perform Fisher zero calculations using ED and tensor network methods. The tensor network results are accurate for most zero points, except for some deviations from the ED results in the relatively larger region of $\beta_{i}$. Since the system is approaching the gapless point, more precise tensor network calculations require a larger $D_{b}$. We find that in the complex $\beta$ plane near $\beta_{r}$ axis, the open line-shaped zeros completely disappear, leaving only zeros that form closed loops. This is similar to the behavior near the QCP in the 1DTFIM.
Figure S4: The Fisher zeros of two different models, obtained from ED and tensor network calculation at $L=16$, are consistent. (a) ANNNI model with $\kappa=0.1$, $g=0.5$ and $D_{b}=30$ (b) XY model with $\gamma=0.01$ and $D_{b}=50$.
The results of the Fisher zeros in these two models indicate that the zero structures exhibited by the 1DTFIM and their relation to quantum criticality have a certain generality. By analyzing the variation of zeros with parameters in these and other models, we can explore the properties of various many-body systems.
II.4 Scaling of the Zeros near QCP
@@
III Comparing the Suzuki solution and CFT at the quantum critical point
In the thermodynamic limit $L\rightarrow\infty$ and at the zero temperature $\beta\rightarrow\infty$, the partition function formula for the Ising critical universality described by Conformal Field Theory (CFT) can be derived from the Suzuki formula. At the critical point $g=g_{c}=1$, the relationship $\epsilon_{k}=2\sin(\pi k/2L)$ holds. Therefore, $\epsilon_{2L}$ in the second term of the Suzuki solution vanishes, and $Z$ can be rewrote as
$Z=$ $\frac{1}{2}\left[e^{2\beta\sum_{k=1}^{L}\sin(\pi k/L)}\prod_{k=1}^{L}\left(1+e^{-4\beta\sin(\pi k/L)}\right)+e^{2\beta\sum_{k=1}^{L}\sin[\pi(k-1/2)/L]}\prod_{k=1}^{L}\left(1+e^{-4\beta\sin[\pi(k-1/2)/L]}\right)\right.$
$\left.+e^{2\beta\sum_{k=1}^{L}\sin[\pi(k-1/2)/L]}\prod_{k=1}^{L}\left(1-e^{-4\beta\sin[\pi(k-1/2)/L]}\right)\right].$
At large $L$, using
$\sum_{k=1}^{L}\sin\frac{\pi k}{L}=\cot\frac{\pi}{2L}=\frac{2L}{\pi}-\frac{\pi}{6L}+O(\frac{1}{L^{3}}),$
and considering that at $k\sim 0$, $\sin(\pi k/L)=\sin[\pi(L-k)/L]\sim\pi k/L$ and at $k\sim L/2$, $e^{-4\beta\sin(\pi k/L)}\sim e^{-4\beta}$ is tiny at large $\beta$, we have
$\prod_{k=1}^{L}\left(1+e^{-4\beta\sin(\pi k/L)}\right)=2\prod_{k=1}^{\infty}\left(1+e^{-4(\beta/L)\pi k)}\right).$
Therefore, the first term in the Suzuki solution is then
$e^{4L^{2}(\beta/L)/\pi}2\tilde{q}^{\frac{1}{12}}\prod_{k=1}^{\infty}(1+\tilde{q}^{k})^{2},$
where $q\equiv e^{-2\pi\beta/L}$ and $\tilde{q}=q^{2}$.
By using
$\sum_{k=1}^{L}\sin\frac{\pi}{L}(k-1/2)=\csc\frac{\pi}{2L}=\frac{2L}{\pi}+\frac{\pi}{12L}+O(\frac{1}{L^{3}}),$
and treating the second and third terms in the Suzuki solution similarly, the full partition function can be expressed as
$Z=e^{4L^{2}(\beta/L)/\pi}\left[2\tilde{q}^{\frac{1}{12}}\prod_{k=1}^{\infty}(1+\tilde{q}^{k})^{2}+\tilde{q}^{-\frac{1}{24}}\prod_{k=1}^{\infty}(1+\tilde{q}^{k-1/2})^{2}+\tilde{q}^{-\frac{1}{24}}\prod_{k=1}^{\infty}(1-\tilde{q}^{k-1/2})^{2}\right].$
In CFT, $\theta(\tau)$ function and Dedekind $\eta(\tau)$ function are usually used, with
$\theta_{2}(\tau)=$ $2q^{\frac{1}{8}}\prod_{k=1}^{\infty}(1-q^{k})(1+q^{k})^{2},$
$\theta_{3}(\tau)=$ $\prod_{k=1}^{\infty}(1-q^{k})(1+q^{k-1/2})^{2},$
$\theta_{4}(\tau)=$ $\prod_{k=1}^{\infty}(1-q^{k})(1+q^{k-1/2})^{2},$
$\eta(\tau)=$ $q^{\frac{1}{24}}\prod_{k=1}^{\infty}(1-q^{k}),$
where $\tau$ is the modular parameter and $q=e^{2\pi i\tau}$. For 1DTFIM, by defining $\tau\equiv i\beta/L$, the partition function is then
$Z=\frac{1}{2}e^{-i4\tau L^{2}/\pi}\left[\frac{\theta_{2}(2\tau)}{\eta(2\tau)}+\frac{\theta_{3}(2\tau)}{\eta(2\tau)}+\frac{\theta_{4}(2\tau)}{\eta(2\tau)}\right],$
which is consistent with the CFT result of the 2D classical Ising model [11], except for the substitution of $\tau$ with $2\tau$. This $2\tau$ term in $Z$ causes the quantum recurrence time $t^{*}$ to be halved compared to the CFT result, so that $t^{*}=4L$ for the 1DTFIM [Fig. S6(a)]. Actually, this does not contradict the result $t^{*}=8L$ in Ref. [12]. In Ref. [12], the Hamiltonian for the 1DTFIM includes a factor of $1/2$, which, after the Jordan-Wigner transformation, results in a form with a single fermionic excitation, thereby matching the result with the Ising CFT. In fermionic representation, the 1DTFIM form in this paper is then corresponds to two fermionic excitations. Since the period depends on the system’s energy scale, doubling the energy exactly halves the period.
Figure S6: The oscillatory behavior of $S(\beta_{r},t)$ at the QCP with different $L$ and $\beta_{r}$. (a) Fine structures of $S$ show the same periodicity $t/L=4$. (b) Example of self similarity of $S$ with $n=3$.
IV Self-similarity of $S$
Self-similarity is a fundamental characteristic of quantum critical points, reflecting the scale-invariant nature of quantum fluctuations at these points. The relationship $S(\beta_{r},t)_{L}\approx S(n\beta_{r},nt)_{nL}$ at $g_{c}=1$ provides a classic example of how systems at different scales can exhibit similar behavior. This property can be used for RG analysis, a powerful tool in studying critical phenomena. Figure S6(b) illustrates the self-similar behavior of $S$ using $n=3$ as an example. The oscillatory behavior of $S$ at different $L$ values not only exhibits the same periodicity but also shows nearly perfect overlap, demonstrating the self-similar nature. The self-similarity implies that the critical exponents, which characterize the behavior of physical quantities near the QCP, are scale-invariant.